OpenStudy (anonymous):

Ms. Chang has 16 students in her class. She wants to send 3 of her students to pick up books for the class. How many combinations of students can she choose? A. 3360 B. 560 C. 250 D. 48 How many combinations of students can she choose?

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

you are being asked to compute \[\binom{16}{3}\] sometimes written as \[_{16}C_3\] and computed via \[\frac{16\times 15\times 14}{3\times 2}\] cancel first, multiply last

OpenStudy (anonymous):

sorry i didnt understand the question Jerk

OpenStudy (anonymous):

haha dont mess with me